Lena29:

The marching band is traveling from Minnesota to Florida to perform in a national competition. A train ticket from Minnesota to Florida cost $150 and takes 15 hours of travel time per student. To rent a bus to make the trip it cost $120 per student and takes 20 hours of travel time per student. The band has $4800 to spend on the trip and can only spend 720 hours of total travel time. What is the maximum number of students who can attend the trip?

Ari1234442:

What are the choises

Mercury:

old question but will respond so this can be closed we can solve this by writing two equations: one representing the total cost, and another representing the total travel time let's let t = # of train tickets and b = # of bus tickets since 1 train ticket is $150, (t) train tickets is $150t same logic with the bus tickets: $120b total cost from tickets: $150t + $120b = $4800

Mercury:

now, for travel time: the train takes 15 hrs of travel time per student, so travel time by train = 15t same logic with the bus: 20b so 15t + 20b = 720 now, your system is: 150t + 120b = 4800 15t + 20b = 720 solve for t and b. then add the t + b to get the total # of students.